OpenStudy (tia_sweetheart17):

simplify the radical. radical 175m^13

OpenStudy (anonymous):

Note that \(\large 175 m^{13} = 7\cdot 25\cdot m\cdot m^{12} = (25m^{12})\cdot 7m = (5m^6)^2\cdot 7m\) Thus, \(\large\sqrt{175m^{13}} = \sqrt{(5m^6)^2\cdot 7m} = \ldots \) Can you finish things off from here? :-)

OpenStudy (tia_sweetheart17):

sure